OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

550=35×3x divided by 35 we get 110/7=3x change of base get ln(110/7)ln(3) 2.5 hours

OpenStudy (campbell_st):

so is the equation \[y = 35 \times 3^x\]

OpenStudy (anonymous):

yea

OpenStudy (campbell_st):

ok... so to graph the equation substitute some values x = 0, 1, 2, 3, 5, 10 which will give you some y values plot them or use some graphing software. you need to graph as you're asked to estimate from the graph... so when you have graphed it look at 550.. the graph is exponential... so looks a bit like 

OpenStudy (anonymous):

i wanted to know am i missing any step in my answer

OpenStudy (anonymous):

yes, it looks like you are missing the step between 110/7=3x and using the change of base, the step where you use log (base 3) on both sides to get \[\log_3 (110/7)=x\]